OpenStudy (anonymous):

Consider the following equation: 4 + 6x = 6x + 4. Explain why the equation has many solutions.

OpenStudy (michele_laino):

here, please subtract 4 from both sides, first

OpenStudy (anonymous):

I don't think they want you to solve it. They want you to explain why the solution has many solutions.

OpenStudy (misty1212):

HI again!!

OpenStudy (anonymous):

Hi love!

OpenStudy (misty1212):

the commutative law says \(a+b=b+a\) always

OpenStudy (michele_laino):

ok! Another methos is the following: please note that the left side is equal to the right side, so your equatiuon is an identity, and an identity is checked for infinite values of x

OpenStudy (misty1212):

therefore \[4+6x=6x+4\] always

OpenStudy (misty1212):

it is the commutative law, but with variables, like \(2+5=5+2\) it is \(4+6x=6x+4\) same thing

OpenStudy (anonymous):

Awesome! How about this one: 2 + 2x = 3 + 2x. explain why the equation does NOT have a solution

OpenStudy (misty1212):

x is a variable, but whatever it is on the left, it is the same number on the right you got a number, you add 2 to it, it has to be different then if you add 3 to it

OpenStudy (michele_laino):

please subtract from both sides 2x

OpenStudy (anonymous):

Thanks!!!!!!

OpenStudy (michele_laino):

thanks!

OpenStudy (misty1212):

like if i have ten dollars and you have ten dollar, i get 2 more, you get 3 more, we cannot have the same amount of money!

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

I'll post another question soon

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]